A vinta is moving up and down periodically, due to waves on the surface of the water. It takes 3.0s for the vinta to travel from its highest to its lowest point, a total distance of 0.70 m. The fisherman on the vinta sees that the wave troughs are spaced 6.0 m apart.


 (a) How fast are the waves traveling?

(b) What is the amplitude of each wave?



1
Expert's answer
2021-07-12T12:17:23-0400

(a) We can find the speed of the waves as follows:


"v=\dfrac{\lambda}{T}=\dfrac{6.0\ m}{2\cdot3.0\ s}=1.0\ \dfrac{m}{s}."


(b) We can find the amplitude of the waves as follows:


"A=\dfrac{0.70\ m}{2}=0.35\ m."